OpenStudy (anonymous):

An executive committee consists of 10 members: 4 men and 6 women. 3 members are selected at random to attend a meeting in Hawaii. The names are drawn from a hat. What is the probability that all 3 selected are men?

OpenStudy (callisto):

P(3 men) = 4/10 x 3/9 x 2/8 = 1/30

OpenStudy (anonymous):

Beat me to it Callisto :P

OpenStudy (callisto):

Sorry wombat..........

OpenStudy (anonymous):

Thank you!

OpenStudy (callisto):

welcome, do you understand how to do it?

OpenStudy (anonymous):

iz all good bro XD

OpenStudy (callisto):

Okay....

OpenStudy (anonymous):

Yes, I knew how to do it prior to this, I just hadn't done it in a few months. I remember now. Thanks again.

OpenStudy (callisto):

Welcome :)